How do I solve this? Set up an equation and solve. The length of a rectangle is twice its width, if the perimeter is 36 meters, what are its demensions? what is the formula for this?
Answer by nerdybill(7384)

How do I solve this? Set up an equation and solve. The length of a rectangle is twice its width, if the perimeter is 36 meters, what are its demensions? what is the formula for this?
.
Let x = width
then
2x = length
.
2(width+length) = perimeter
2(x+2x) = 36
2(3x) = 36
(3x) = 18
x = 6 meters (width)
.
length:
2x = 2(6) = 12 meters
.
dimensions: 6 meters by 12 meters